How do you solve the following equation:
2x squared minus 5x equals 12

Answer by edjones(8007)

2x^2-5x=12
2x^2-5x-12=0
2*-12=-24; The two factors of -24 whose sum is -5 are -8 and +3.
2x^2+3x-8x-12=0
x(2x+3)-4(2x+3)=0
(x-4)(2x+3)=0
x=4,
2x=-3, x=-3/2
.
Ed